in determining whether a regression slope is significant, first you calculate t. what is the value of t if b= 8 and the standard error of b =8?

socoilogist use different statistical teats partly based on the level if measurement of the variables they are examining. what is the most common situation in which you'd use crosstab and chi-squared?

we test for a difference in family size between cuba and italy by randomly selecting 100 cubans and 100 italians and asking them how many family memebers reside in their households. what is the degrees of freedom if we test for a difference in these two means using a t-distribution?
Answer #1

1] NULL HYPOTHESIS H0:

ALTERNATIVE HYPOTHESIS Ha:

Test statistic t= coefficient of b1/SE(b1)

t= 8/8

t= 1.0

So test statistic t=1.0

2] THE MOST COMMON SITUATION WHEN WE USE CROSS TAB AND CHI SQARE WHEN WE WANT TO TEST ASSOCIATION BETWEEN BETWEEN TWO RANDOM VARIABLES.

3] DEGREES OF FREEDOM= 100+100-2= 200-2=198